OpenStudy (anonymous):

Find the simplified form of the expression. Give your answer in scientific notation. please explain (3 * 10^8)(9 * 10^7)

OpenStudy (anonymous):

multiply the numbers out front add the exponents

OpenStudy (anonymous):

are u familiar with the doorkeeper method

OpenStudy (anonymous):

i.e. \[3\times 9=27\\ 10^8\times 10^7=10^{8+7}=10^{15}\]

OpenStudy (anonymous):

no i am not

OpenStudy (anonymous):

ohhh ok i get it haha thanx. :)

OpenStudy (anonymous):

@Squeeks1713 be careful with your answer

OpenStudy (anonymous):

i was asking @Squeeks1713 actually

OpenStudy (anonymous):

no im not, sorry

OpenStudy (anonymous):

it is not \[27\times 10^{15}\] because that is not in scientific notation it is the right number, but not in scientific notation because \(27>10\)

OpenStudy (anonymous):

yeah i know, i was just saying to never heard of it, figured i would learn something

OpenStudy (anonymous):

right i got that. haha thank you so much :)